What is his marginal rate of substitution between X =3 and Y=4? If A=1, a = 0.4 and b =
0.5.
1
Expert's answer
2020-07-10T10:59:09-0400

If the utility function is "U = AX^aY^b" , then marginal utility functions for X and Y are:

"MU(X) = U'(X) = A\u00d7a\u00d7X^{a - 1}\u00d7Y^b."

"MU(Y) = U'(Y) = A\u00d7b\u00d7X^a\u00d7Y^{b - 1}."

"MRS(XY) = MU(X)/MU(Y) = \frac{a×X^{a - 1}×Y^b} {b×X^a×Y^{b - 1}} = \frac{aY}{bX} = \frac{0.4\times4}{0.5\times3} = 1.067."